OpenStudy (anonymous):

simplfiy x^2-9x+14/ x^2+4x-12

OpenStudy (anonymous):

\[\frac{ x^2-9x+14 }{ x^2+4x-12 }\]

OpenStudy (anonymous):

answers \[\frac{ x-7 }{ x-2 } where X \neq2\] \[\frac{ x-2 }{ x-6 } whersX \neq 6\] \[\frac{ x+7 }{ x-6 } whereX \neq 6\] \[\frac{ x-7 }{ x+6 } where X \neq 6,2\]

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

See if you can tell what comes next. If you mess up, no worries.